Part A When 1.550 g of liquid hexane (C6H14) undergoes combustion in a bomb calorimeter, the...

Part A

When 1.550 g of liquid hexane (C6H14) undergoes combustion in a bomb calorimeter, the temperature rises from 25.87∘C to 38.13∘C. Find ΔErxn for the reaction in kJ/mol hexane. The heat capacity of the bomb calorimeter, determined in a separate experiment, is 5.73 kJ/∘C.

Express your answer in kilojoules per mole to three significant figures.

ΔErxn

ΔErxn

of hexane =

nothing

kJ/mol

SubmitRequest Answer

Part B

The combustion of toluene has a ΔErxn of –3.91×103 kJ/mol. When 1.55 g of toluene (C7H8) undergoes combustion in a bomb calorimeter, the temperature rises from 23.12∘C to 37.57∘C. Find the heat capacity of the bomb calorimeter.

Express the heat capacity in kilojoules per degree Celsius to three significant figures.

Ccal

Ccal

=

nothing

  kJ/∘C  
Answer #1

Part A:

ΔErxn = -C * delta T

= - 5.73*(38.13 - 25.87)

= -70.25 KJ

number of moles = mass/molar mass

= 1.55 / (6 * 12 + 14 * 1)

= 0.018 mol

ΔErxn = -70.25 KJ / 0.018 mol

= -3.90 * 10^3 KJ/mol

Part B:

number of moles = mass/molar mass

= 1.55/ (7*12 + 8*1)

= 0.017 mol


ΔErxn = ΔErxn per mol * number of moles

   = -3.91 * 10^3 * 0.017

= -65.875 KJ

ΔErxn = C * T

-65.875 = -C * (37.57 - 23.12)

C = 4.56 KJ/oC